ANSWER: We begin by looking at the summary statistcs, a dotplot, and a boxplot of the median teachers pay across all states. Summary statistics: n Mean Variance Std. dev. Std. err. Median Range Min Max Q1 Q3 Column Teacher Pay ($1000 51 35.890196 38.762502 6.2259539 0.87180792 792 35 24 26.3 50.3 31.5 40.5 30 45 35 40 5o Teacher Pay ($1000 so 30 5 45 Teacher Pay($1000 Nationwide, the median median teacher pay (not a mistake to repeat the word median twice here) is $35,000. The minimum median teacher pay is in the state of The maximum median teacher pay is Although one could argue that the distribution is slightly skewed , the distribution is mostly symmetric, with an IQR of % of the total range of in the state of to the which is just - Connected to Microsoft Exch
The mean median teacher pay is,with a standard deviation of With a relatives small sample size of Nit is hard to say too much more abou this distribution. In the data, there are two ways we can look at the question of regional differences. We can either look at the salaries based on census region or more simply by breaking down all of the states into two categories: coastal states versus central states. If we look at the census regions first, we get the following summary statistics, dotplots and boxplots. Summary statistics for Teacher Pay ($1000) Group by: Region Region n . Mean Variance . Std. dev. . Std. err. Median . Ranges Min Ma. Qī。Q3: S 39.88 9.287 3.047458 1.3628646 3 7.1 37.7 44.8 37.8 40.9 6.48 2.5455844 1.2727922 3. 5.4 27.7 33.1 29.5 33.1 2 16.1 48.1 461 48.1 ENC 31.3 4 ESC 47.9 3963 6.3073) 43759 0 0233125 62 1%382 30: 10 3 47.366667 1.2133333 1.1015141 0.63595947 MA 8 32.025 6.4707143 2.5437599 0.89935493 31.256.8 29.4 36.2 30.1 MTN 6 40.066667 40.210667 6.3411881 2.5887792 39.25 17.4 32.9 50.3 35.8 42.9 NE PAC 5 41.22 29.032 5.3881351 2.4096473 39.6 13.8 35.8 49.6 38 43.1 SA 7 31.785714 15.481429 3.9346447 1.4871559 32.4 10.6 26.3 36.9 27 35.1 5.2 26.8 32 27.6 30.65 WNC 4 29.125 4.7425 2.1777282 1.0888641 28.85 WSC Region wsc.
uddies WindowHelp s TakeHomeExam1Spring 2017.pdf (page 2 of 7) Q Search WSC PAC MTN MA ESC ENC 50 45 35 30 Teacher Pay ($1000) wsc WNC MTN 50 40 Teacher Pay ($1000) 45 35 30
Clearly, the median teachers pay varies considerably from census region to census region. Identify the five regions with the lowest median median teachers pay. What do these five regions have in common that may explain why they pay teachers general less than states in other census regions? Identify the four regions with the highest median median teachers pay What do these four regions have in common that may explain why they pay teachers general more than states in other census regions?
ddies Window Help sTakeHomeExam1Spring2017.pdf (page 4 of 7) Q Search Finally, we look again at median teachers salary by state, but instead of breaking the states down into census regions, we will use the simpler designation of coastal or central states. Coastal means that the state has a coast with the Atlantic or Pacific Ocean, or the Gulf of Mexico. Otherwise, a state is considered central. Answer the question: Is there a noticeable difference in the median teachers pays in coastal versus central states? Summary statistics for Teacher Pay ($1000): Group by: Coast/Central Coast/Centrale n Mean Variance Std. dev.. Std. err..Median Range Min Max Q1 03 27 33.566667 18.288462 4.2765011 0.82301302 33.1 18.5 26.3 44.8 30.6 36.3 24 38.504167 50.124764 7.0798845 1.4451754 38.8 23.5 26.8 50.3 32.45 43.4 Central Coast .s Coas. CentraleI 30 50 40 45 35 Teacher Pay ($1000) Coast, Central Coast Connected to Microsoft Ex
TakeHomeExam1Spring2017.pdf (page 4 of 7) Q Search Coost 30 35 40 45 50 Teacher Pay ($1000 Coast Central 30 35 40 45 SO Teacher Pay ($1000
Wed May 29 6:50 PM Allie O. t 93%-, Buddies Window Help " D TakeHomeExam1Spring2017.pdf (page 5 of 7) 2. (/35 pts) Using Question 1 as a model, answer the following question. Keep your lives simple - only break down the states by coastal of central states. Answer this question thoroughly on your own sheets of paper and include appropriate displays of data. Choose displays that illustrate the points you are trying to make, and refer to them in your written narrative. The written narrative should not be longer than two paragraphs. Are there noticeable differences among coastal and central states in the percent of students who take the SAT Exam? If so, describe the differences.
/25 pts) Pose a question similar to the question in Question 2, focusing your question 3. the SAT Verbal scores, the SAT Math scores, or the Percent with No High School Diploma, broken down by coastal and central states. Answer this question thoroughly on your own sheets of paper and include appropriate displays of data. Choose displays that illustrate the points you are trying to make, and refer to them in your written narrative. The written narrative should not be longer than two paragraphs
Answer #1

1)
median = 35000
minimum = 26300
maximum = 50300
skewed to the right (As mean > median)
IQR = Q3- Q1 = 40500 - 31500 = 900

Total range = 50300-26300
= 2400
hence % = 900/2400 = 37.5%


mean is 35890
standard deviation is 6225.95
N = 51